At 0 degrees C the density of liquid water is 0.9999gml-1 what the volume of 1.000g of liquid water
1
Expert's answer
2020-02-13T06:15:32-0500

p=m/V.

V=m/p= 1/0.9999 = 1.00010001 ml.